# Downsizing Diffusion Models for Cardinality Estimation

### v1

@@ Abstract.
Inspired by the performance of score-based diffusion models (Song et al., 2021b) in estimating complex text, video, and image distributions with thousands of dimensions, we introduce Accelerated Diffusion Cardest (ADC), the first joint distribution (Wang et al., 2021) cardinality estimator based on a downsized diffusion model.
To calculate the pointwise density value of data distributions, ADC’s density estimator $p_{\theta}(x)$ uses a formula (Song et al., 2021a) that evaluates log-likelihood by integrating the score function $\nabla\log p_{t}(x)$, a gradient mapping which ADC has learned to efficiently approximate using its lightweight score estimator, $s_{\theta}(x,t)$. To answer ranged queries, ADC’s selectivity estimator first predicts their selectivity using a Gaussian Mixture Model (GMM) $q(x)$, then uses importance sampling Monte Carlo to correct its predictions with more accurate pointwise density values $\{p_{\theta}(x_{i})\}_{i=1}^{n}$ calculated by the density estimator. ADC+ further trains a decision tree to identify the high-volume, high-selectivity queries that $q(x)$ alone can predict very accurately, in which case it skips the correction phase to prevent Monte Carlo from adding more variance. Doing so lowers median Q-error and cuts per-query latency by 25%, making ADC+ usually twice as fast as Naru (Yang et al., 2019), arguably the state-of-the-art joint distribution cardinality estimator.
Numerical experiments  [^1] Open-source code for ADC shall be released at https://github.com/XinheMu/ADC-Replication- in Jan. 2026 upon formal submission using well-established benchmarks (Asuncion et al., 2007) (Wang et al., 2021) show that on all real-world datasets tested, ADC+ is capable of rivaling Naru and outperforming MSCN (Kipf et al., 2018), DeepDB (Hilprecht et al., 2020), LW-Tree, and LW-NN (Dutt et al., 2019) using around 66% their storage space, being at least $3\times$ as accurate as MSCN on 95th and 99th percentile error [^2] We define ”accuracy” as Q-error1-1 as 1 is Q-error’s optimal value. Furthermore, on a synthetic dataset where attributes exhibit complex, multilateral correlations, ADC and ADC+ are considerably robust while almost every other learned model suffered significant accuracy declines. In this case, ADC+ performs better than any other tested model, being $10\times$ as accurate as Naru (Yang et al., 2019) on 95th and 99th percentile error.
Cardinality Estimation, Diffusion Model, Score Matching
copyright: none
@@ 1. Introduction
Score-based Diffusion Models have presented the deep learning community with a very, if not the most, attractive way of acquiring highly local information about a complex high-dimensional probability distribution (Chen et al., 2023). These models work by training a neural network to approximate the score (gradient of its log function) of an evolving probability density function, which, starting out as the distribution to be learned, is gradually turned into Gaussian noise through a diffusion process. Using the learned score function, score-based diffusion models can then reverse the diffusion process by solving a backward diffusion stochastic differential equation (SDE), and, in doing so, sample from (Song et al., 2021b)(Ramesh et al., 2022)(Dhariwal and Nichol, 2021) or evaluate (Song et al., 2021b)(Song et al., 2021a) the underlying distribution at the user’s request.
Given their success, one may naturally wonder whether cardinality estimation, a task crucial for the ideal performance of query optimizers, can also benefit from diffusion models. Indeed, several learned cardinality estimation models, such as Naru(Yang et al., 2019) and DQM-D(Hasan et al., 2020), are already performing cardinality estimation by approximating an underlying probability density function at individual points, then numerically integrating it across the queried region using various Monte Carlo schemes(Yang et al., 2019)(Lepage, 1978). However, simply grabbing a popular score-based diffusion model off the shelf probably won’t work due to the three reasons below.
First, while conventional models used for image generation excel at approximating distributions with thousands of dimensions, their sampling process often take up to tens of seconds even with the help of advanced TPUs(Ho et al., 2020) and time-saving techniques like consistency models(Song et al., 2023) or DPM Solver(Lu et al., 2022), which is often longer than the time it costs to actually execute the query.
Second, existing works on diffusion models have never touched on the problem of integrating the probability density function across a given region, a task useless for image generation, yet crucial for cardinality estimation.
Most importantly, image generation and cardinality estimation require different magnitudes of precision: while making it five times more likely to generate a specific image gives a generative model its unique art style, predicting five thousand tuples to exist in a region where only one thousand actually does gives a query optimizer an optimization disaster.
In this paper, we aim to overcome the aforementioned limitations by presenting Accelerated Diffusion Cardest(ADC), a “downsized” diffusion model much more suited for cardinality estimation, structured according to the framework in Figure 1.
@@
    Downsizing Score Matching. We find that for distributions with moderate dimensionality, the inherent structure and properties of diffusion processes can be used to construct much more lightweight score matching networks that achieve the same level of precision. That is, score matching is best done using the data prediction model when $t$ is large, and a modified noise prediction model when $t$ is small. We modify traditional noise prediction models by constructing a neural network with three output modules, normalized using the functions $\frac{1}{\sigma^{2}(t)},\frac{1}{\sigma(t)}$, and $1$, respectively. This modification is proven to significantly enhance precision and, to the best of our knowledge, has not been proposed in previous works.
@@
    Simplifying Likelihood Calculation. Existing works studying diffusion models (Song et al., 2021a) provide a formula for calculating the pointwise log-density of a probability distribution by integrating its score function across time and space. We convert this formula into a fast, well grounded algorithm which combines Quasi Monte Carlo (for integration across space) with adaptive stepsize Midpoint Rule (for integration across time), enabling accurate estimations in less than $1\text{\,}\mathrm{ms}$ per point.
@@
    GMM for Cardinality Estimation. In order to estimate the selectivity of ranged queries, we turn to the importance sampling algorithm $\int_{V}p(x)dx=\int_{V}q(x)dx\mathrm{E}_{q(x|x\in V)}\frac{p(x)}{q(x)}$, where the problem boils down to constructing a suitable predictor $q$. Here, we propose training a Gaussian Mixture Model (GMM) for $q$, as conditional sampling, pointwise density calculation, and integration across hyperrectangles all prove very easy for GMMs. Numerical experiments find that GMMs not only cooperate very well with diffusion models, but also prove to be efficient standalone estimators. This is especially true for queries with large selectivity or great volumes, an advantage we harnessed in ADC+ by training a decision tree to identify queries ”easy” enough for the GMM $q(x)$ to handle independently.
  4. (4)
    Experimental Evaluation Highlights. We test ADC and ADC+ against five state-of-the-art learned models (DeepDB(Hilprecht et al., 2020), Naru(Yang et al., 2019), LW-NN, LW-Tree(Dutt et al., 2019), MSCN(Kipf et al., 2018)), strictly adhering to well-established benchmarks for dataset choice and workload design(Wang et al., 2021). We find that on real-world datasets, ADC+ is comparable to Naru and almost always beats MSCN, DeepDB, LW-Tree, and LW-NN, using less than 360KB of storage space and with a latency at most half that of Naru’s. Further experiments using synthetic datasets show that ADC works best on relations where strong synergistic interactions exist between attributes, quantifiable by a large and positive interaction information (i.e. one can, and can only, significantly narrow down the range of one attribute by knowing several others). In this case, no other learned model performs as well as ADC+, and Naru even loses to ADC+ on 95%th and 99%th Q-error by a more than 10 $\times$ margin.
2. Background, Notations, and Related Work
@@ 2.1. Cardinality Estimation
The goal of cardinality estimation is to rapidly and accurately estimate the number of rows returned by a given query without actually executing it. This essay, in particular, studies a more specific and highly influential subcase presented by (Heimel et al., 2015) and (Dutt et al., 2019) (implied in their problem formulation and experiment dataset choice, though LW-Tree and LW-NN versions supporting categorical attributes are later developed (Wang et al., 2021)), in which the underlying database is a single relation $\mathrm{R}$ with continuous attributes $\{A_{1},...,A_{n}\}$, and the query $\mathrm{Q}$ is of the form
SELECT count(*) FROM $\mathrm{R}$
WHERE $a_{k_{1}}<A_{k_{1}}<b_{k_{1}}$
AND $a_{k_{2}}<A_{k_{2}}<b_{k_{2}}$
AND …
AND $a_{k_{i}}<A_{k_{i}}<b_{k_{i}}$,
i.e. the queried region is always a hyper-rectangle.
Current approaches to cardinality estimation can be very roughly categorized into learned and non-learned methods (Wang et al., 2021). Traditional non-learned methods, such as small-scale sampling, multidimensional histograms (Bruno et al., 2001), Bayesnet (Chow and Liu, 1968), and kernel density estimators (Heimel et al., 2015), are usually more time-efficient and relatively robust to constant updates (Wang et al., 2021), but much less accurate than learned models. Therefore, they are of more use when the data distribution is simple, when the database is frequently updated, or when the precision requirements are not strict.
Learned models, on the other hand, usually conduct cardinality estimation in one of the following two ways (Wang et al., 2021). Regression models, such as MSCN (Kipf et al., 2018) and LW-tree (Dutt et al., 2019), learn from already-conducted queries by conveying them into feature vectors, then constructing a regression model to match these vectors with the true cardinality of their corresponding queries. Joint distribution models, meanwhile, assume that data points are distributed according to an unknown probability density function, and thus turn cardinality estimation into a numerical integration problem: namely, calculating the density function at individual points, then summing them up using a predefined integration scheme. For example, estimators like Naru (Yang et al., 2019) and DQM-D (Hasan et al., 2020) calculate the joint distribution function at each point by factorizing it into conditional distributions
@@
then sum them up using the Monte Carlo scheme
$\int_{V}p(x)dx=\int_{V}g(x)dx\cdot\mathrm{E}_{g(x|x\in V)}\frac{p(x)}{g(x)},$
with the predictor, $g(x)$, constructed in multiple stages using sequential sampling (Naru), VEGAS (Lepage, 1978) (DQM-D), or other sampling algorithms.
Given that estimating and sampling from complex, high-dimensional distributions is exactly what score-based diffusion models do best, as is proven by their groundbreaking success in image, text, and video generation (Oko et al., 2023), it is no surprise that in the field of joint distribution estimators, diffusion models have the full potential to walk under the spotlight.
@@ 2.2. Score-Based Diffusion Models
Score-based diffusion models are the generative models that made AI drawing and video generation a reality, thanks to pioneering work by Song, Dickstein and Kingma (Song et al., 2021b), Song and Dhariwal (Song et al., 2023), Ho, Jain, and Abbeel (Ho et al., 2020), Leobacher and Pillichshammer (Dhariwal and Nichol, 2021), and many others. For the sake of convenience, all mathematical notations related to diffusion models shall be documented in Appendix A. The meaning of each notation will also be explained upon its first appearance.
As mentioned before, given a set of data points satisfying an unknown distribution $p_{0}$ in the sample space $\mathbb{R}^{d}$, score-based diffusion models attempt to learn and sample from $p_{0}$ in two steps: the Forward Process and the Backward Process(Song et al., 2021b).
The Forward Process gradually injects white noise into the sample points to create Gaussian kernels around them, slowly turning $p_{0}$ into a smooth Gaussian distribution using the SDE
$dx=-\alpha(t)xdt+\beta(t)dw,t\in[0,T],$
where $dw$ refers to the standard Brownian motion.
Two most common choices for $\alpha$ and $\beta$ are $\alpha(t)=0,\beta(t)=1$ (the VE scheme) and $\alpha(t)=1,\beta(t)=\sqrt{2}$ (the VP scheme, which is used in our model)(Song et al., 2021b). However, works including (Lee et al., 2022) and (Zhang et al., 2024) show that we can, in theory, choose any positive function for $\alpha$ and $\beta$ without fundamentally changing the model. In fact, for each choice of $\alpha$ and $\beta$, it would always hold that  [^3] See Appendix B for proof

### v2

@@ Abstract.
Learned cardinality estimation requires accurate model designs to capture the local characteristics of probability distributions. However, existing models may fail to accurately capture complex, multilateral dependencies between attributes. Diffusion models, meanwhile, can succeed in estimating image distributions with thousands of dimensions, making them promising candidates, but their heavy weight and high latency prohibit effective implementation. We seek to make diffusion models more lightweight by introducing Accelerated Diffusion Cardest (ADC), the first ”downsized” diffusion model framework for efficient, high-precision cardinality estimation. ADC utilizes a hybrid architecture that integrates a Gaussian Mixture-Bayesnet selectivity estimator with a score-based density estimator to perform precise Monte Carlo integration. Addressing the issue of prohibitive inference latencies common in large generative models, we provide theoretical advancements concerning the asymptotic behavior of score functions as time $t$ approaches zero and convergence rate estimates as $t$ increases, enabling the adaptation of score-based diffusion models to the moderate dimensionalities and stringent latency requirements of database systems.
We also introduce ADC+, an optimized variant that dynamically identifies queries with high volume and selectivity, bypassing complex density evaluations in these cases to reduce variance and latency. Through experiments conducted against five learned estimators, including the state-of-the-art Naru, we demonstrate that ADC and ADC+ offer superior robustness when handling datasets with multilateral dependencies, which cannot be effectively summarized using pairwise or triple-wise correlations. In fact, ADC+ is 10 times more accurate than Naru on such datasets. Additionally, ADC+ achieves competitive accuracy comparable to Naru across all tested datasets while maintaining latency half that of Naru’s and requiring minimal storage (¡350KB) on most datasets.
@@ 1. Introduction
Score-based Diffusion Models have presented the deep learning community with a very, if not the most, attractive way of acquiring highly local information about a complex high-dimensional probability distribution (reftwentyseven). These models work by training a neural network to approximate the score (gradient of its log function) of an evolving probability density function, which, starting out as the distribution to be learned, is gradually turned into Gaussian noise through a diffusion process. Using the learned score function, score-based diffusion models can then reverse the diffusion process by solving a backward diffusion stochastic differential equation (SDE), and, in doing so, sample from (refthirteen)(reftwelve)(refone) or evaluate (reftwo)(refone) the underlying distribution at the user’s request.
Given their success, one may naturally wonder whether cardinality estimation, a task crucial for the ideal performance of query optimizers, can also benefit from diffusion models. Indeed, several learned cardinality estimation models, such as Naru(refsix) and DQM-D(refseven), are already performing cardinality estimation by approximating an underlying probability density function at individual points, then numerically integrating it across the queried region using various Monte Carlo schemes(refeleven)(refsix). However, simply grabbing a popular score-based diffusion model off the shelf probably won’t work due to the three reasons below.
First, while conventional models used for image generation excel at approximating distributions with thousands of dimensions, the sampling process often takes up to tens of seconds, even with the help of advanced TPUs(refeight) and time-saving techniques like consistency models(reffour) or DPM Solver(reften), which is often longer than the time it costs to actually execute the query.
Second, image generation and cardinality estimation require different magnitudes of precision: while making it five times more likely to generate a specific image gives a generative model its unique art style, predicting 5000 tuples to exist in a region where only 1000 actually do gives a query optimizer an optimization disaster.
Most importantly, existing works on diffusion models have never touched on the problem of integrating the probability density function across a given region, a task useless for image generation, yet crucial for cardinality estimation.
We aim to overcome the aforementioned limitations by presenting Accelerated Diffusion Cardest (ADC) and ADC+  [^1] See https://github.com/XinheMu/ADC-Replication- for open source code of ADC, downsized diffusion models suited for cardinality estimation with their structure described in Figure 1, where
  • ADC answers ranged queries via a prediction-sampling-correction Monte Carlo algorithm, and
  • ADC+ takes a shortcut to skip sampling and correction when answering high volume, high selectivity queries, for which the prediction is already accurate enough.
@@
    Downsizing Diffusion Models. We present the first application of diffusion models to cardinality estimation, providing theoretical guidance that results in more expressive networks and better parameter choices under moderate dimensionalities. We overcome existing challenges by presenting two major findings regarding the diffusion process and the score function. That is, we give asymptotic estimates for the score function as $t\to 0$ around regions with different continuity conditions, helping us mitigate the problem of score blowup by introducing normalization coefficients that grow at different speeds, and we give lower bounds for the speed at which $\nabla\log p_{t}(x)$ converges to $\frac{x}{\sigma^{2}(t)}$ as $t$ grows large, which proves faster than what is generally assumed (reftwentyseven), helping us choose the integration time-step and $T$, the perturbation upper bound. Using these results, one can design diffusion models that are more accurate and efficient in the database scenario, utilizing their relative freedom from the curse of dimensionality.
@@
    Constructing ADC/ADC+. Using such findings, we design ADC, a diffusion-based algorithm that estimates query cardinality using three modules: a Score Estimator $s_{\theta}(x,t)$ to match the score function $\nabla\log p_{t}(x)$, a Density Estimator $p_{\theta}(x)$ to estimate pointwise densities using an integration formula for likelihood evaluation (reftwo), and a Gaussian Mixture Model(GMM)-Bayesnet hybrid Selectivity Estimator $q(x)$ to be the predictor in the Monte Carlo formula
    (1) $\int_{V}p(x)dx=\int_{V}q(x)dx\cdot\mathrm{E}_{q(x|x\in V)}\frac{p(x)}{q(x)}$
    for ranged queries. ADC+ further uses features ”predicted selectivity $\int_{V}q(x)dx$” and ”query volume” to identify queries for which the variance added by $\mathrm{E}_{q(x|x\in V)}\frac{p(x)}{q(x)}$ exceeds the bias it corrects, in which case it trusts the selectivity estimator to directly output $\int_{V}q(x)dx$, resulting in lower median Q-error and latency compared to ADC.
@@
    Experimental Evaluation. We test ADC and ADC+ against five learned estimators (DeepDB(refthirtythree), Naru(refsix), LW-NN, LW-Tree(refeighteen), MSCN(refseventeen)), adhering to well-established benchmarks for dataset choice and workload design(reffive). We find that ADC and ADC+ are able to learn complex, multilateral dependencies between attributes that cannot be captured by pairwise or triple-wise correlations. In fact, on a synthetic dataset designed to magnify the above trait, ADC+ performs best in the models tested, being 10 times more accurate  [^2] We define ”accuracy” as the difference between Q-error and 1, as 1 is Q-error’s optimal value than Naru on 95%th and 99%th Q-error. ADC+ also performs well on real-world datasets, rivaling Naru in accuracy and almost always beating all other models, using less than 350KB of storage space and with a latency of less than 10ms per query for the majority of datasets.
2. Background and Related Work
@@ 2.1. Cardinality Estimation
2.1.1. Problem Formulation
Consider a relation $R$ with attribute domains $\{A_{1},...,A_{d}\}$ and a query $Q$, which can be viewed as a subset $\Omega\subseteq A_{1}\times A_{2}\times...\times A_{d}$, cardinality estimation requires us to estimate the query’s cardinality, i.e., $|\Omega\cap R|$, without conducting a full search of the relation $R$.
More specifically, existing research on cardinality estimation, such as (reffourteen) and (refeighteen), often consider an important subproblem in which $\{A_{i}\}_{i=1}^{d}$ are all bounded intervals of $\mathbb{R}$ and $\Omega=\prod_{i=1}^{d}B_{i}$, with $B_{i}\subseteq A_{i}$ a subinterval of $A_{i}$ (possibly $A_{i}$ itself). This is also the subproblem we study throughout the paper.
2.1.2. Current Methods
Current approaches to cardinality estimation can be very roughly categorized into learned and non-learned methods (reffive). Traditional non-learned methods, such as small-scale sampling, multidimensional histograms (reffifteen), Bayesnet (refsixteen), and kernel density estimators (reffourteen), are usually more time-efficient and relatively robust to constant updates (reffive), but less accurate than learned models. Thus, they are of more use when the data distribution is simple, when the database is frequently updated, or when the precision requirements are not strict.
Learned models, on the other hand, usually conduct cardinality estimation in one of two ways (reffive). Regression models, such as MSCN (refseventeen), LW-tree, and LW-NN (refeighteen), learn from labeled queries by converting them into feature vectors, then constructing a regression model to match these vectors with the true cardinality of corresponding queries. Besides producing estimation results more accurate than traditional models, these models also have the advantage of taking relatively little time to train, and can make an estimate as quickly as many non-learned methods.
Joint distribution models, meanwhile, assume that data points are distributed according to a probability density function $p(x)$, and thus turn cardinality estimation into a numerical integration problem: that is, calculating $p(x)$ at individual points, then summing them up using a predefined integration scheme. For example, estimators like Naru (refsix) and DQM-D (refseven) calculate the joint distribution function by factorizing it into conditional distributions
@@ 2.1.2. Current Methods
then sum them up using the Monte Carlo formula (1) with the predictor, $q(x)$, constructed via sequential sampling (Naru), VEGAS (refeleven) (DQM-D), or other algorithms. While joint distribution models cost more time to train and infer, they are known to produce some of the most accurate cardinality estimators.
2.1.3. Why Diffusion Models?
The high accuracy of joint distribution models invites us to dig deeper into this specific category. However, current joint distribution models all share a subtle but common limitation: conventional methods of constructing the estimator $p$, be it autogression (Naru, DQM-D) or sum-product networks (DeepDB (refthirtythree)), tend to treat a tuple as a collection of correlated, but ultimately distinct, set of attributes rather than a single unified point, risking declines in accuracy if different attributes exhibit complex, multilateral correlations (eg. lying a short distance from several disjoint submanifolds of high curvature) that cannot be efficiently captured by pairwise or even triple-wise correlation alone. While Naru, arguably the state-of-the-art joint distribution estimator, attempts to handle this by factorizing autoregression models in the most expressive way possible (i.e., considering all prior attributes when predicting a new one), it thus experiences a high latency, and, as shown in our experiments, still struggles to capture enough details when the distribution grows too complex.
A diffusion model, on the other hand, generates an image by solving a unified reverse diffusion equation, rather than predicting the values of independent pixel clusters. What’s more, their accuracy and expressiveness are repeatedly proven by their success in the generation of images and videos, whose dimensionality often soars in the realm of thousands. Therefore, they bring the potential of overcoming the previous limitation, taking the achievements of joint distribution models one step further.
@@ 2.2. Score-Based Diffusion Models
2.2.1. Introduction and Notations
Score-based diffusion models are the generative models that made AI drawing and video generation a reality, thanks to pioneering work by Song, Dickstein and Kingma (refone), Song and Dhariwal (reffour), Ho, Jain, and Abbeel (refeight), Leobacher and Pillichshammer (refthirteen), and many others. For clarity, we present a summary of notations in Table 1. Each notation will also be explained upon its first appearance.
Table 1. Mathematical Notations
Notation Definition
$d$ Dimensionality of the dataset
$w$ Standard Brownian motion
$\hat{w}$ Standard Wiener process in reverse time
$\varphi_{\sigma^{2}}$ Distribution density function of $\mathcal{N}(0,\sigma^{2})$
$\tau_{k}$ Scaling operator $(\tau_{k}\circ g)(x)=k^{d}g(kx)$
$\Phi$ Cumulative density function of $\mathcal{N}(0,1)$
$p_{0}(x)$ The original data distribution we seek to reconstruct
$k_{\alpha,\beta}(t)$ $e^{\int_{0}^{t}\alpha(s)ds}$, abbreviated as $k(t)$
$\sigma^{2}_{\alpha,\beta}(t)$ Function satisfying $\frac{d}{dt}(\sigma^{2}(t))=\beta(t)-2\alpha(t)\sigma^{2}(t)$, $\sigma^{2}(0)=0$, abbreviated as $\sigma^{2}(t)$
$p_{t,\alpha,\beta}(x)$ Distribution derived by perturbing $p_{0}$ using the diffusion SDE $dx=-\alpha(t)dt+\beta(t)dw$ from time 0 to $t$, abbreviated as $p_{t}(x)$
$p_{0t,\alpha,\beta}(x|x_{0})$ Distribution derived by perturbing $\delta_{x_{0}}$, a distribution concentrated on $x_{0}$, using the above SDE from time 0 to $t$, abbreviated as $p_{0t}(x|x_{0})$.
$q(x)$ Distribution reconstructed by the GMM predictor.
$q_{t,\alpha,\beta}(x)$ Distribution derived by perturbing $q_{0}=q$ using the above SDE from time 0 to $t$, abbreviated as $q_{t}(x)$.
$\epsilon$ Early stopping time implemented upon training $s_{\theta}(x,t)$, see 3.3 for why it is needed.
$\tilde{s}_{\theta}(x,t)$ $s_{\theta}(x,t+\epsilon)$
2.2.2. Forward and Backward process
Given a set of data points satisfying an unknown distribution $p_{0}$ in the sample space $\mathbb{R}^{d}$, score-based diffusion models attempt to learn and sample from $p_{0}$ in two steps: the Forward Process and the Backward Process(refone).
The Forward Process gradually injects white noise into the sample points to create Gaussian kernels around them, slowly turning $p_{0}$ into a smooth Gaussian distribution using the forward Stochastic Differential Equation (SDE)
(2) $dx=-\alpha(t)xdt+\beta(t)dw,t\in[0,T],$
where $w$ refers to the standard Brownian motion.
Two most common choices for $\alpha$ and $\beta$ are $\alpha(t)=0,\beta(t)=1$ (the VE scheme) and $\alpha(t)=1,\beta(t)=\sqrt{2}$ (the VP scheme)(refone). However, works including (refthirtytwo) and (reftwentyfour) show that we can, in theory, choose any positive function for $\alpha$ and $\beta$ without fundamentally changing the model, as is formally stated by the theorem below. Therefore, we shall assume the VP scheme in all later discussions unless otherwise specified.
Theorem 2.1.
 [^3] See Appendix A for proof of all theorems
Let $p_{0}$ be a probabilistic distribution and $\{p_{t}|t\in[0,T]\}$ be a family of distributions derived by perturbing $p_{0}$ using the SDE (2) from time 0 to time $T$. Let $k(t)=e^{\int_{0}^{t}\alpha(s)ds}$, and $\sigma^{2}(t)$ be the solution of the initial value Ordinary Differential Equation (ODE)
${d\sigma^{2}}/{dt}$ $=\beta(t)-2\alpha(t)\sigma^{2}(t)$
$\sigma^{2}(0)$ $=0,$
then, for each choice of $\alpha$ and $\beta$, it would hold that